In 1755 Bret Bighound made a bank account and put 10$ in it on January 1st 1755. the bank account has an annual interest rate of 5%. since then the interest has been COMPOUNDED quarterly. If you are correct in calculating the money as of December 31st 2000, you will be withdrawing the money out. My question is : Before you withdraw it, how much money will be in the account? I need to show all work. Thank you.
Answer by stanbon(75887)

In 1755 Bret Bighound made a bank account and put 10$ in it on January 1st 1755. the bank account has an annual interest rate of 5%. since then the interest has been COMPOUNDED quarterly. If you are correct in calculating the money as of December 31st 2000, you will be withdrawing the money out. My question is : Before you withdraw it, how much money will be in the account? I need to show all work.
--------
2000-1755 = 245
t = 246
------------------
A(t) = P(1+(r/n))^(nt)
Ans:: 10(1+(0.05/1))^(1*246) = 10*163142.67 = $1,631,427
-----------
Cheers,
Stan H.
------------